MICROWAVE SPECTRUM AND STRUCTURE OF HBr-SO_{3} AND THE QUADRUPOLE COUPLING CON-STANT OF Ar-^{33}SO_{3}

\underline{DENISE FIACCO} , M.E. OTT, B.HARMS, W.A. BURNS, M.CANAGARATNA, and K.R. LEOPOLD, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455.

We report the microwave spectrum and structure of HBr-SO_{3} as determined by pulsed nozzle Fourier transform microwave spectroscopy. Like its previously studied HF and HCl analogs, the complex exhibits a symmetric top spectrum and has a structure in which the halogen approaches the sulfur along the C_{3} axis of the SO_{3}. The S–Br distance is 3.24 A and the HBr subunit forms an angle of approximately 73 deg. (or 107 deg.) with the a-axis of the complex. The structures of the three HX-SO_{3} complexes are compared. We also report a measurement of the ^{33}S nuclear quadrupole coupling constant in Ar-^{33}SO_{3}. The value obtained (-15.296(14) MHz) is compared with a recent theoretical calculation of the monomer value and discussed in terms of the angular oscillations of the SO_{3} unit.